If a football kicked from the ground moves with a velocity of 20 MS-1.making 30 degree with the horizontal, find its vertical displacement and horizontal displacement after one second

Grade:8

1 Answers

Eshan
5 years ago
Dear student,

Component of velocity along horizontal=u_x=ucos30^{\circ}=10\sqrt{3}m/s
Component of velocity along vertical=u_y=usin30^{\circ}=10m/s

Hence horizontal displacement in 1 second=s_x=u_xt=10\sqrt{3}m
Vertical displacement in 1 second=s_y=u_yt-\dfrac{1}{2}gt^2=(10)(1)-5(1^2)=5m